OpenStudy (anonymous):

It took 1700 J of work to stretch a spring from its natural length of 2m to a length of 5m. Find the spring's force constant.

OpenStudy (nottim):

You're looking for "x"?

OpenStudy (anonymous):

Well constant is K.. I'm just confused by this whole chapter so far.

OpenStudy (anonymous):

I know it uses Hooke's Law, which is F = K * X

OpenStudy (nottim):

X is 5-2. Use the Work to find the F (there a formula? I forgot) You're looking for "K" I guess. Ignore the already given constant.

OpenStudy (anonymous):

1700 = K - 3?

OpenStudy (anonymous):

I mean K * 3

OpenStudy (nottim):

I don't think so. Force is not work. And yes, 3k works for the right side.

OpenStudy (nottim):

Work equals "Force times distance" right? W=f*d.

OpenStudy (nottim):

So, I guess "d" is probably your "x". Sub in 1700 for W.

OpenStudy (anonymous):

F = -kx 1700/3 = -k(5-2) 1700/9 = -k k = -1700/9

OpenStudy (anonymous):

1700=3F

OpenStudy (nottim):

@ Chris: 1700J isn't force...Is it?

OpenStudy (nottim):

Yup. Rearrange 1700=3f so that 1700/3=f. There's your force.

OpenStudy (anonymous):

W = Fs 1700 = F(3) F = 1700/3

OpenStudy (anonymous):

None of those have worked for answers.. :/

OpenStudy (anonymous):

what are you trying to find

OpenStudy (nottim):

Well that's only a part of it...

OpenStudy (nottim):

He's looking for "K".

OpenStudy (nottim):

But you basically have your "f" and "X". You already tried it with f as 566.666... and x as 3?

OpenStudy (anonymous):

yeah I tried +/ 1700/9 and 1700/3

OpenStudy (anonymous):

try this \[F = \Delta K\] 1700 = (1/2)k(3)^2 - 0 (2*1700)/(9) = k

OpenStudy (nottim):

Did you get 188.9?

OpenStudy (anonymous):

@Chris, that was correct!

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

so how did you get that equation?

OpenStudy (nottim):

Aww. I fail at physics.

OpenStudy (anonymous):

Yeah this is calculus 2, but for engineers.. Even though I'm a computer science major.

OpenStudy (anonymous):

F = -kx <===== this is for when x's natural position is at 0 You have to integrate it if the natural position is elsewhere \[\int\limits_{0}^{x} F dx = \int\limits_{0}^{x}kx dx = (1/2) kx^2 |0 \to x = (1/2)kx^2 - 0\]